OpenStudy (anonymous):

can someone help me solve this problem: square root of 2t and square root of 2t

OpenStudy (anonymous):

\[\sqrt{2t} \sqrt{2t}\]

OpenStudy (anonymous):

can someone please help me solve the problem that I just posted?

OpenStudy (anonymous):

\[\sqrt{4t^{2}}\]

OpenStudy (anonymous):

which is 2t

OpenStudy (anonymous):

oh, I had 2 for the answer

OpenStudy (anonymous):

always get whwat is under radical when you square square roots

OpenStudy (anonymous):

awesome! Thank you so much!

OpenStudy (anonymous):

welcome...medal please : )

OpenStudy (anonymous):

how do I give you a medal?

OpenStudy (anonymous):

look on end of line by my name